# Lie symmetry group methods for differential equations

### v10

@@ 2 Differential equations and their symmetry group
$\Phi_{\varepsilon}(x,u):\quad\tilde{x}(x,u;\varepsilon)=\frac{x}{1-\varepsilon x},\quad\tilde{u}(x,u;\varepsilon)=\frac{u}{1-\varepsilon x},$
@@ 3.1 Invariant differentiation
$[y_{1}^{2}\partial_{y_{1}}+3y_{1}y_{2}\partial_{y_{2}}+(3y_{2}^{2}+4y_{1}y_{3})\partial_{y_{3}}]I=0$
@@ 4 Reduction of order for ordinary differential equations
This means that Eq. (52) can be integrated by quadrature (separation of variables). This equation involves interesting types of ODEs like Abel’s equation of the second type or Riccati equation.
@@
We construct the most general second order ODE invariant under the rotation group $\operatorname{SO}(3,\mathbb{R})$. The Lie algebra $\operatorname{so}(3,\mathbb{R})$ acting on the surface of the unit sphere $\mathbb{S}^{2}=\left\{(x,y):x\in[0,2\pi),y\in(0,\pi)\right\}\subset\mathbb{R}^{3}$ is realized by the vector fields
@@
A first integral is immediately obtained as $\tau=\csc yy_{1}$. The second independent integral is found by integrating the linear ODE
$\frac{dy_{2}}{dy}-2\cot yy_{2}=-(1+2\tau^{2})$
@@
Finally, a single invariant equation depending on an arbitrary parameter $\nu$ is found from the relation $\zeta(y,y_{1},y_{2})=\nu=\text{constant}$. In terms of the spherical variables $(x,y)$ it can be written in the form
$y_{2}-2(\cot y)y_{1}^{2}-\sin y\cos y=\nu\sin^{2}y(1+\csc^{2}y)^{3/2}.$ (72)
@@
In order to solve this equation, we can use the algorithmic method of first integrals based on symmetries suggested in [18]. We first need the characteristic functions of the vector fields $\mathbf{v}_{1},\mathbf{v}_{2},\mathbf{v}_{3}$
@@
$c_{1}\sin x+c_{2}\cos x+\cot y=0.$
@@
The first one is the conservation of energy in which $\mathscr{I}_{1}=H=L-y_{1}L_{y_{1}}$ is the Hamiltonian. Of course, only two of them are functionally independent. Indeed, we have the relation
@@
which is in fact the (general) solution with the parameter $y_{1}$ eliminated. We can put $c_{1}=1$ by dividing by $c_{1}$ and redefining $c_{2}$ and $c_{3}$.
@@
$\mathbf{v}_{2}=a(x,y)\partial_{x}+b(x,y)\partial_{y},\quad\mathbf{v}_{2}=A(x,y)\partial_{x}+B(x,y)\partial_{y}$
@@
$\gamma(\sin\sigma)=0,\quad-1-\gamma^{2}-\gamma^{\prime}(\cos\sigma)+\gamma^{\prime}\sigma^{\prime}(\sin\sigma)=0.$
@@
All locally primitive Lie algebras of vector fields in $\mathbb{R}^{2}$ isomorphic to are equivalent to the Lie algebra spanned by the vector fields
@@
So we find a family of invariant equations $I_{4}=I_{3}+2I_{2}^{-1/2}=c$, depending on single parameter, or in terms of $(x,y)$ coordinates
@@
The second order invariant equation with the same symmetry $\mathbf{v}_{1}=\xi\partial_{x}+\eta y\partial_{y}$ of the previous example can be expressed in terms of the second order invariants; $r,w$ as defined in (50) and
@@
If $\xi=x^{2}$ and $\eta=x$ are chosen, $\mathbf{v}_{1}$ generates an inversional group and this equation simplifies to
$y^{\prime\prime}=x^{-3}H(r,w),\quad r=\frac{y}{x},\quad w=xy^{\prime}-y.$
@@
Theorem 4.8 guarantees that integration is completed using two quadratures by reductions in the order of $\mathbf{v}_{1}$ (an ideal of $\mathfrak{g}$) first and then $\mathbf{v}_{2}$. We remark that apart from the special case $y^{\prime\prime}=Kx^{-5}y^{2}$, which is obtained for $n=2$ from (91), there are only two other values of the exponent $m$ in the equation $y^{\prime\prime}=Kx^{m}y^{2}$ for which the symmetry algebra is a two-dimensional (rank-two) non-abelian one. They are $m=-15/7$ and $m=-20/7$ (see [51] for classification details). For all values $m$ with $m\notin\left\{-5,-15/7,-20/7\right\}$, the symmetry algebra is one-dimensional and generated by the scaling symmetry $\mathbf{v}=x\partial_{x}-(m+2)y\partial_{y}$.
The case $n=-3$ gives the Ermakov–Pinney equation
@@
$y=\pm\sqrt{A+2Bx+Cx^{2}},\quad AC-B^{2}=K,$
which arises in many applications. This equation admits the $\operatorname{sl}(2,\mathbb{R})$ algebra as the symmetry algebra with the basis
@@
if the condition $03+\delta+m-k(m+1)=$ 0 is satisfied. The case when $\delta=0$ ($m=(k-3)/(1-k)$) appeared in [12] as one of the reductions of a radially symmetric nonlinear porous-medium equation. A suitable basis of the algebra in the case $k\neq 1,2$ is
@@
If we identify (100) with (97) we find $p=c$, $q=1$ and $m=2$, $K=1$ and the integrability condition imposes a constraint $c=\pm 5/\sqrt{6}$ on the wave speed. The corresponding solution is given in terms of Weierstrass function by $\tilde{y}=\wp(1/\sqrt{6}(\tilde{x}+a);0,g_{3})$ (see page 177).
@@
The expression on the left-hand side of (121) is called the Schwarzian derivative of $y$ with respect to $x$ and is denoted by the symbol $\left\{y,x\right\}$. It is invariant under the Möbius transformation in $y$: $\left\{(ay+b)/(cy+d);x\right\}=\left\{y,x\right\}$, $ad-bc=1$ (a unique differential invariant of order $\leq 3$ of the algebra (122)). A two-dimensional solvable subalgebra $\mathfrak{g}_{0}$ generated by $\left\{\mathbf{v}_{1},\mathbf{v}_{2}\right\}$ can be used to reduce the equation to one of first order. The second-order differential invariants of $\mathfrak{g}_{0}$ are $x$ and $w=y_{2}/y_{1}$, in terms of which the reduced equation becomes
@@
There are only two types of third order ODE admitting 6-dimensional symmetry groups: $(1+y_{1}^{2})y_{3}=3y_{1}y_{2}^{2}$ and $\left\{y,x\right\}=0$. The maximal symmetry algebra for a third order ODE is seven-dimensional and is attained if and only if it is equivalent to the linear equation $y_{3}=0$ up to a linear point (equivalence) transformation.
@@ 4.1 Group classification problem
Note that the above result also contains the subcase $k=1$ (but not $k=-1$). By change of basis $\mathbf{v}_{1}\to-\mathbf{v}_{1}$, $\mathbf{v}_{2}\to\mathbf{v}_{2}$ and $\mathbf{v}_{1}\to k\mathbf{v}_{1}$, $\mathbf{v}_{2}\to\mathbf{v}_{2}$ the commutation relation takes the standard form for a two-dimensional non-abelian algebra: $[\mathbf{v}_{1},\mathbf{v}_{2}]=\mathbf{v}_{2}$.
@@
$\begin{split}&\xi^{\prime\prime}-2\eta^{\prime}+3(\alpha f-\alpha^{\prime\prime}y)=0,\\ &\xi(x,y)f_{x}+\eta(x,y)f_{y}=(\eta-2\xi^{\prime})f+\alpha^{\prime\prime\prime}y^{2}+\eta^{\prime\prime}y+\beta^{\prime\prime}.\end{split}$ (152)
@@
Hence, if $\alpha=0$, then $\xi^{\prime\prime}-2\eta^{\prime}=0$ and $f(x,y)$ is determined from
$\xi(x)f_{x}+(\eta(x)y+\beta(x))f_{y}=(\eta-2\xi^{\prime})f+\eta^{\prime\prime}y+\beta^{\prime\prime}.$
@@
If we introduce the canonical variables for the one-dimensional subgroups of $\operatorname{SL}(2,\mathbb{R})$ symmetry group corresponding to the vector fields $\mathbf{v}_{4},\mathbf{v}_{5},\mathbf{v}_{4}+\mathbf{v}_{6}$
$r=\frac{y}{\sqrt{\xi}},\quad s=\int\frac{dx}{\xi(x)},\quad\xi\in\left\{y_{1}^{2},y_{1}y_{2},y_{2}^{2}\right\},$
then the transformed ODE becomes a constant-coefficient one in the form
$\frac{d^{2}r}{ds^{2}}+\varepsilon W^{2}r=0,\quad\varepsilon\in\left\{0,-1,1\right\}.$ (156)
Here, $W$ is the Wronskian of $y_{1},y_{2}$ (a constant).
@@
allowing the two-dimensional non-abelian symmetry algebra spanned by the vector fields

### v11

@@ 2 Differential equations and their symmetry group
$\Phi_{\varepsilon}(x,u):\quad\tilde{x}(x,u;\varepsilon)=\frac{x}{1-\varepsilon x},\quad\tilde{u}(x,u;\varepsilon)=\frac{u}{1-\varepsilon x},$ (22)
@@ 3.1 Invariant differentiation
$\left[y_{1}^{2}\partial_{y_{1}}+3y_{1}y_{2}\partial_{y_{2}}+(3y_{2}^{2}+4y_{1}y_{3})\partial_{y_{3}}\right]I=0$
@@ 4 Reduction of order for ordinary differential equations
This means that Eq. (53) can be integrated by quadrature (separation of variables). This equation involves interesting types of ODEs like Abel equation of the second type or Riccati equation.
@@
We construct the most general second order ODE invariant under the rotation group $\operatorname{SO}(3,\mathbb{R})$. The Lie algebra $\operatorname{so}(3,\mathbb{R})$ acting on the surface of the unit sphere $\mathbb{S}^{2}=\left\{(x,y):x\in[0,2\pi),y\in(0,\pi)\right\}\subset\mathbb{R}^{3}$ (See Remark 4.6) is realized by the vector fields
@@
A first integral is immediately obtained as $\tau=(\csc y)y_{1}$. The second independent integral is found by integrating the linear ODE
$\frac{dy_{2}}{dy}-2(\cot y)y_{2}=-(1+2\tau^{2})$
@@
Finally, a single invariant equation depending on an arbitrary parameter $\nu$ is found from the relation $\zeta(y,y_{1},y_{2})=\nu=\text{constant}$. In terms of the spherical angle variables (azimuth and polar angles) $(x,y)$ it can be written in the form
$y_{2}-2(\cot y)y_{1}^{2}-\sin y\cos y=\nu\sin^{2}y\left(1+(\csc^{2}y)y_{1}^{2}\right)^{3/2}.$ (73)
@@
In order to solve this equation, we can use the algorithmic method of first integrals based on symmetries suggested in [18] (see [51] for an alternative integration strategy). We first need the characteristic functions of the vector fields $\mathbf{v}_{1},\mathbf{v}_{2},\mathbf{v}_{3}$
@@
$c_{1}\sin x+c_{2}\cos x+\cot y=0$ (79)
or in terms of new parameters $\alpha,\beta$ as
$\cot y=\alpha\cos(x+\beta).$ (80)
This is the equation of great circle (intersection of the plane passing through the origin with the unit sphere).
@@
The first one is the conservation of energy in which $\mathscr{I}_{1}=-H=L-y_{1}L_{y_{1}}$ is the Hamiltonian. Of course, only two of them are functionally independent. Indeed, we have the relation
@@
which is in fact the general solution (without any integration) with the parameter $y_{1}$ eliminated. If $c_{1}\neq 0$, we can put $c_{1}=1$ by dividing by $c_{1}$ and redefining $c_{2}$ and $c_{3}$.
Alternatively, the first integral $\mathscr{I}_{1}$, which is invariant under $\partial_{x}$,
$c\sin^{4}y=y_{1}^{2}+\sin^{2}y,\quad c=c_{1}^{-2}$
can be written as
$[(\cot y)^{\prime}]^{2}+(\cot y)^{2}=c-1\equiv\alpha,$
which integrates to the solution (80).
@@
$\mathbf{v}_{2}=a(x,y)\partial_{x}+b(x,y)\partial_{y},\quad\mathbf{v}_{3}=A(x,y)\partial_{x}+B(x,y)\partial_{y}$
@@
$\gamma(\sin\sigma)=0,\quad-(1+\gamma^{2})-\gamma^{\prime}(\cos\sigma)+\gamma^{\prime}\sigma^{\prime}(\sin\sigma)=0.$
@@
As a result, we can state that any second order ordinary differential equation admitting a rank-two realization of $\operatorname{so}(3,\mathbb{R})$ algebra is equivalent to (73) by point transformations.
@@
All locally primitive Lie algebras of vector fields in $\mathbb{R}^{2}$ isomorphic to $\operatorname{so}(3,\mathbb{R})$ are equivalent to the Lie algebra spanned by the vector fields
@@
So we find a family of invariant equations $I_{4}=I_{3}+2I_{2}^{-1/2}=c$, depending on a single parameter, or in terms of $(x,y)$ coordinates
@@
The second order invariant equation with the same symmetry $\mathbf{v}_{1}=\xi\partial_{x}+\eta y\partial_{y}$ of Example (4.2) can be expressed in terms of the second order invariants; $r,w$ as defined in (51) and
@@
The particular choice $\eta=\xi^{\prime}/2$ leads to the simpler equation
$y^{\prime\prime}+q(x)y=\xi^{-3/2}H(r,w),\quad q(x)=-\frac{(\sqrt{\xi})^{\prime\prime}}{\sqrt{\xi}},$ (94)
where the invariant variables $r,w$ have the form
$r=\frac{y}{\sqrt{\xi}},\quad w=\frac{1}{\sqrt{\xi}}\left(\xi y^{\prime}-\frac{\xi^{\prime}}{2}y\right).$
If $\xi=x^{2}$ is chosen, $\mathbf{v}_{1}$ generates an inversional group and this equation simplifies to
$y^{\prime\prime}=x^{-3}H(r,w),\quad r=\frac{y}{x},\quad w=xy^{\prime}-y.$ (95)
@@
Theorem 4.8 guarantees that integration is completed using two quadratures by reductions in the order of $\mathbf{v}_{1}$ (an ideal of $\mathfrak{g}$) first and then $\mathbf{v}_{2}$. We remark that apart from the special case
$y^{\prime\prime}=Kx^{-5}y^{2},$ (97)
which is obtained for $n=2$ from (96) , there are only two other values of the exponent $m$ in the equation $y^{\prime\prime}=Kx^{m}y^{2}$ for which the symmetry algebra is a two-dimensional (rank-two) nonabelian one. They are $m=-15/7$ and $m=-20/7$ (see [51] for classification details). For all values $m$ with $m\notin\left\{-5,-15/7,-20/7\right\}$, the symmetry algebra is one-dimensional and generated by the scaling symmetry $\mathbf{v}=x\partial_{x}-(m+2)y\partial_{y}$.
Eq. (97) is invariant under the two-dimensional symmetry algebra
$\mathbf{v}_{1}=x^{2}\partial_{x}+xy\partial_{y},\quad\mathbf{v}_{2}=x\partial_{x}+3y\partial_{y}.$ (98)
We can integrate (97) using differential invariants $I=y/x$ and $J=xy^{\prime}-y$ of $\mathbf{v}_{1}$. The reduced first order equation is obtained by invariant differentiation
$\frac{dJ}{dI}=\frac{x^{3}y^{\prime\prime}}{xy^{\prime}-y}=K\frac{x^{-2}y^{2}}{xy^{\prime}-y}=K\frac{I^{2}}{J},$ (99)
whose integration gives
$J^{2}-\frac{2K}{3}I^{3}=c_{1}$
or
$(xy^{\prime}-y)^{2}-\frac{2K}{3}x^{-3}y^{3}=c_{1}.$ (100)
Note that this equation is invariant under the scaling group $(x,y)\to(\lambda x,\lambda^{3}y)$, $\lambda>0$ corresponding to $\mathbf{v}_{2}$. In terms of the normal coordinates $r=y/x$, $s=-1/x$ of $\mathbf{v}_{1}$ it has the form
$r^{\prime}(s)^{2}=\frac{2K}{3}r^{3}+c_{1}.$
For $c_{1}\neq 0$, the general solution can be expressed in terms of elliptic functions. It is elementary, when $c_{1}=0$, which is given by $r=(1/(\sqrt{6})s+c_{2})^{-2}$ and in terms of $(x,y)$ variables
$y(x)=\frac{6x^{3}}{(\sqrt{K}+\sqrt{6}c_{2}x)^{2}}.$ (101)
Moreover, the particular case $n=-3$ of (96) gives the celebrated Ermakov–Pinney equation
@@
$y=\pm\sqrt{A+2Bx+Cx^{2}},\quad AC-B^{2}=K.$
This equation admits the $\operatorname{sl}(2,\mathbb{R})$ algebra as the symmetry algebra with the basis
@@
if the condition $03+\delta+m-k(m+1)=$ 0 is satisfied.
The case when $\delta=0$ ($m=(k-3)/(1-k)$) appeared in [12] as one of the reductions of a radially symmetric nonlinear porous-medium equation. A suitable basis of the algebra in the case $k\neq 1,2$ is
@@
If we identify (110) with (107) we find $p=c$, $q=1$ and $m=2$, $K=1$ and the integrability condition imposes a constraint $c=\pm 5/\sqrt{6}$ on the wave speed. The corresponding solution is given in terms of Weierstrass function by $\tilde{y}=\wp(1/\sqrt{6}(\tilde{x}+a);0,g_{3})$ (see page 199).
@@
The expression on the left-hand side of (131) is called the Schwarzian derivative of $y$ with respect to $x$ and is denoted by the symbol $\left\{y,x\right\}$. It is invariant under the Möbius transformation in $y$:
$\left\{\frac{ay+b}{cy+d};x\right\}=\left\{y,x\right\},\quad ad-bc=1$
(a unique differential invariant of order $\leq 3$ of the algebra (132)). A two-dimensional solvable subalgebra $\mathfrak{g}_{0}$ generated by $\left\{\mathbf{v}_{1},\mathbf{v}_{2}\right\}$ can be used to reduce the equation to one of first order. The second-order differential invariants of $\mathfrak{g}_{0}$ are $x$ and $w=y_{2}/y_{1}$, in terms of which the reduced equation becomes
@@
There are only two types of third order ODE admitting 6-dimensional symmetry groups:
$(1+y_{1}^{2})y_{3}=3y_{1}y_{2}^{2}\quad\mbox{and}\quad\left\{y,x\right\}=0.$
The maximal symmetry algebra for a third order ODE is seven-dimensional and is attained if and only if it is equivalent to the linear equation $y_{3}=0$ up to a linear point (equivalence) transformation.
@@ 4.1 Group classification problem
Note that the above result also contains the subcase $k=1$ (but not $k=-1$). By change of basis $\mathbf{v}_{1}\to-\mathbf{v}_{1}$, $\mathbf{v}_{2}\to\mathbf{v}_{2}$ and $\mathbf{v}_{1}\to k\mathbf{v}_{1}$, $\mathbf{v}_{2}\to\mathbf{v}_{2}$ the commutation relation takes the standard form for a two-dimensional nonabelian algebra: $[\mathbf{v}_{1},\mathbf{v}_{2}]=\mathbf{v}_{2}$.
@@
$\begin{split}&\xi^{\prime\prime}-2\eta^{\prime}+3(\alpha f-\alpha^{\prime\prime}y)=0,\\ &\xi f_{x}+\eta f_{y}=(\eta-2\xi^{\prime})f+\alpha^{\prime\prime\prime}y^{2}+\eta^{\prime\prime}y+\beta^{\prime\prime}.\end{split}$ (162)
@@
Hence, if $\alpha=0$, then $\xi^{\prime\prime}-2\eta^{\prime}=0$ and $\eta=\xi^{\prime}/2+c_{0}$. We have a single classification relation between $f$, $\xi$ and $\beta$.
$\xi(x)f_{x}+(\eta(x)y+\beta(x))f_{y}=(\eta-2\xi^{\prime})f+\eta^{\prime\prime}y+\beta^{\prime\prime}.$ (164)
For given $\xi$, $\beta$ and $c_{0}$, the first order linear PDE (164) can be integrated using method of characteristics to find the most general form of $f$. For $c_{0}=0$, $f(x,y)$ is given by
[…]
where $F$ ($F^{\prime\prime}\neq 0$) is an arbitrary function of its argument. We note that for $\xi=x^{2}$, $\beta=0$ we recover (95) with the arbitrary function $H$ independent of the differential invariant $w$.
Imposing a further symmetry on the equation (165) one can generate a second order differential equation invariant under a two-parameter group of symmetries with the arbitrary function $F$ restricted. As an example, we require invariance under the scaling group $(x,y)\to(\lambda x,\lambda^{k}y)$, $\lambda>0$ and pick $\xi,\beta$ for $c_{0}=0$ such that the generators
$\mathbf{v}_{1}=\xi\partial_{x}+(\frac{\xi^{\prime}}{2}y+\beta)\partial_{y},\quad\mathbf{v}_{2}=x\partial_{x}+ky\partial_{y}$ (166)
form a two-dimensional Lie algebra with commutator
$[\mathbf{v}_{1},\mathbf{v}_{2}]=(\xi-x\xi^{\prime})\partial_{x}+\left(k\beta-x\beta^{\prime}-\frac{\xi^{\prime\prime}}{2}xy\right)\partial_{y}=-\mathbf{v}_{1},$
which implies
$x\xi^{\prime}-2\xi=0,\quad(k+1)\beta-x\beta^{\prime}+\frac{1}{2}(\xi^{\prime}-x\xi^{\prime\prime})y=0$
with solutions $\xi=x^{2}$, $\beta=\delta x^{k+1}$ so we find the generators
$\mathbf{v}_{1}=x^{2}\partial_{x}+(xy+\delta x^{k+1})\partial_{y},\quad\mathbf{v}_{2}=x\partial_{x}+ky\partial_{y}.$
With these choices of $\xi,\beta$, invariance under $\mathbf{v}_{2}$ will restrict the arbitrary $F$ to
$F(\sigma)=F_{0}\sigma^{(k+1)/(k-1)},\quad\sigma=\frac{y}{x}-\frac{\delta}{k-1}x^{k-1},\quad k\neq 1.$ (167)
In conclusion, the corresponding invariant differential equation becomes
$y^{\prime\prime}=\delta kx^{k-2}+F_{0}x^{-3}\sigma^{(k+1)/(k-1)}$ (168)
and can be integrated by two quadratures in the order of $\mathbf{v}_{1},\mathbf{v}_{2}$. The special case $\delta=0$ of (168) agrees with the Emden–Fowler equation (96) for $n=(k+1)/(k-1)$ and in particular with (97) for $k=2$.
When $k=1$,
$F(\sigma)=F_{0}e^{-(2/\delta)\sigma},\quad\sigma=\frac{y}{x}-\delta\log x,\quad\delta\neq 0$ (169)
the equation assumes the form
$y^{\prime\prime}=x^{-1}\left[\delta+F_{0}e^{-(2/\delta)(y/x)}\right],$ (170)
invariant under a rank-2 Lie algebra with generators
$\mathbf{v}_{1}=x^{2}\partial_{x}+x(\delta x+y)\partial_{y},\quad\mathbf{v}_{2}=x\partial_{x}+y\partial_{y}.$
Using the normal coordinates $r=\sigma$ and $s=-1/x$ of $\mathbf{v}_{1}$ satisfying $\mathbf{v}_{1}(\sigma)=0$, $\mathbf{v}_{1}(s)=1$ we can write (170) as
$\frac{d^{2}\sigma}{ds^{2}}=F_{0}e^{-(2/\delta)\sigma},$
which is certainly invariant under $\tilde{\mathbf{v}}_{1}=\partial_{s}$ and $\tilde{\mathbf{v}}_{2}=s\partial_{s}+\delta\partial_{\sigma}$.
We can form an equivalent realization of the Lie algebra spanned by the elements (166) satisfying the commutation relation
$[\mathbf{v}_{1},\mathbf{v}_{2}]=(\xi-x\xi^{\prime})\partial_{x}+\left(k\beta-x\beta^{\prime}-\frac{\xi^{\prime\prime}}{2}xy\right)\partial_{y}=(1-m)\mathbf{v}_{1}.$ (171)
In this case, we have
$\xi=x^{m},\quad\beta=\delta x^{k+m-1}$
for which Eq. (165) remains remains invariant under $\mathbf{v}_{2}$ when $F$ is limited to a power function of $\sigma$
$F=F_{0}\sigma^{\nu},\quad\sigma=x^{-m/2}\left(y+\frac{2\delta x^{k}}{m-2k}\right),\quad\nu=\frac{4-3m-2k}{m-2k},\quad m\neq 2k.$ (172)
The corresponding invariant equation is significantly simplified for $\nu=2$ or $m=2/5(k+2)$ ($k\neq 1/2$). In this case, there are two values of the pairs $(m,k)$, that are worth examining.
First we choose $(m,k)=(6/7,1/7)$ so that $\nu=2$, $\beta=\delta$ (constant). After some calculation, we find the invariant equation as
$y^{\prime\prime}=F_{0}x^{-15/7}y^{2}+\frac{(343F_{0}\delta-12)}{196x^{2}}(7\delta x^{1/7}+4y).$ (173)
This equation is further simplified if we choose $\delta=12/(343F_{0})$
$y^{\prime\prime}=F_{0}x^{-15/7}y^{2}$ (174)
admitting the symmetry algebra
$\mathbf{v}_{1}=x^{6/7}\partial_{x}+(\frac{3}{7}x^{-1/7}y+\frac{12}{343F_{0}})\partial_{y},\quad\mathbf{v}_{2}=x\partial_{x}+\frac{1}{7}y\partial_{y}$
with commutation relation $[\mathbf{v}_{1},\mathbf{v}_{2}]=1/7\mathbf{v}_{1}$.
In a similar way, the choice $(m,k)=(8/7,6/7)$ so that $\nu=2$, $\beta=\delta x$ leads to the invariant equation
$y^{\prime\prime}=F_{0}x^{-20/7}y^{2}+\frac{(343F_{0}\delta+12)}{196x^{2}}(7\delta x^{6/7}-4y).$ (175)
Again the choice $\delta=-12/(343F_{0})$ gives the invariant equation
$y^{\prime\prime}=F_{0}x^{-20/7}y^{2}$ (176)
admitting the symmetry algebra
$\mathbf{v}_{1}=x^{8/7}\partial_{x}+(\frac{4}{7}x^{1/7}y-\frac{12}{343F_{0}}x)\partial_{y},\quad\mathbf{v}_{2}=x\partial_{x}+\frac{6}{7}y\partial_{y}$
with commutation relation $[\mathbf{v}_{1},\mathbf{v}_{2}]=-1/7\mathbf{v}_{1}$. We recall that all possible forms of the equation $y^{\prime\prime}=F_{0}x^{n}y^{2}$ invariant under a two-dimensional nonabelian symmetry algebra are attained only for three values of the exponents $n\in\left\{-5,-15/7,-20/7\right\}$ (see page 97).
In the Abelian case $[\mathbf{v}_{1},\mathbf{v}_{2}]=0$ ($\nu=1$), we find $\xi=x$, $\beta=\delta x^{k}$ and this leads to a linear equation.
On the other hand, if $f$ has some specific form, this linear PDE allows us to find all possible symmetries. For example, if $f(x,y)=q(x)y^{2}$, then we find that $\xi$, $\beta$ should satisfy
$\xi^{\prime\prime\prime}=4q\beta,\quad\beta^{\prime\prime}=0,\quad 2\xi q^{\prime}+(5\xi^{\prime}+2c_{0})q=0.$
Once $\xi$, $\beta$ are solved form the first two equations, the last one serves as a compatibility condition. For the choice $q=e^{x}$, it can be shown that there is only one symmetry given by $\mathbf{v}=\partial_{x}-y\partial_{y}$ which reduces the corresponding equation to an Abel equation of second type. For another special case $q=Kx^{m}$ see the discussion on page 97.
@@
allowing the two-dimensional nonabelian symmetry algebra spanned by the vector fields